(1 point) Use the substitution method to solve the system -x + y = 0, 4x – 3y = -3. Your answer is
Answer #1

Given system of equations

-x + y = 0

4x - 3y = -3

From first equation

-x + y = 0 we get

y = x

Substitute this y = x in 4x - 3y = -3 we get

4x - 3(x) = -3

4x - 3x = -3

x = -3

Substitute this x = -3 in y = x we get

y = -3

Therefore

x = -3 and y = -3